# Improving Alignment and Robustness with Circuit Breakers

### v2

@@ Abstract
AI systems can take harmful actions and are highly vulnerable to adversarial attacks. We present an approach, inspired by recent advances in representation engineering, that interrupts the models as they respond with harmful outputs with “circuit breakers.” Existing techniques aimed at improving alignment, such as refusal training, are often bypassed. Techniques such as adversarial training try to plug these holes by countering specific attacks. As an alternative to refusal training and adversarial training, circuit-breaking directly controls the representations that are responsible for harmful outputs in the first place. Our technique can be applied to both text-only and multimodal language models to prevent the generation of harmful outputs without sacrificing utility—even in the presence of powerful unseen attacks. Notably, while adversarial robustness in standalone image recognition remains an open challenge, circuit breakers allow the larger multimodal system to reliably withstand image “hijacks” that aim to produce harmful content. Finally, we extend our approach to AI agents, demonstrating considerable reductions in the rate of harmful actions when they are under attack. Our approach represents a significant step forward in the development of reliable safeguards to harmful behavior and adversarial attacks. Code is available at github.com/blackswan-ai/circuit-breakers.
Work done while at Black Swan AI. Work done in advising capacity to Black Swan AI.
Correspondence to: andy@blackswan.ai
@@ 3 Circuit Breaking with Representation Engineering
In this section, we introduce a novel approach aimed at mitigating the generation of harmful outputs in neural networks by inducing a new type of phenomenon called “circuit-breaking.” This phenomenon can be elicited using a family of techniques designed to remap model representations related to harmful processes, redirecting them towards incoherent or refusal representations. This process is reminiscent of “short-circuiting,” where harmful representations are “shorted” and intercepted by circuit breakers. The core objective of this method is to robustly prevent the model from producing harmful or undesirable behaviors.
@@ 4.1 Large Language Models
    Input Embedding Attack: This white-box attack operates in the embedding space by optimizing a set of input embeddings directly instead of using hard tokens, with the objective of eliciting an affirmative assistant response.
@@ 4.2 Multimodal Models
Results.  Figure 3 demonstrates that for multimodal models, our circuit-breaking technique RR is also able to make a model significantly more robust while preserving model capabilities. Especially when subject to white-box PGD Attack, RR achieves reduction of $84\%$ in the compliance rate compared to the original model and $85\%$ compared to the safety prompt. Meanwhile, performance on MMMU and LLaVA-Wild remains within $0.5\%$ of the original model’s, as opposed to the safety prompt which causes a decrease of $3.3\%$ on LLaVA-Wild. This demonstrates that despite the ongoing challenge of achieving adversarial robustness in standalone image recognition, circuit breakers enable the larger multimodal system to reliably counter image “hijacks” [6] intended to elicit harmful outputs.
@@ 5 Limitations and Conclusion
We are thankful to Steven Basart, Stephen Casper, David Dalrymple, and Xander Davies for providing valuable feedback on the paper.

### v3

@@ Abstract
AI systems can take harmful actions and are highly vulnerable to adversarial attacks. We present an approach, inspired by recent advances in representation engineering, that interrupts the models as they respond with harmful outputs with “circuit breakers.” Existing techniques aimed at improving alignment, such as refusal training, are often bypassed. Techniques such as adversarial training try to plug these holes by countering specific attacks. As an alternative to refusal training and adversarial training, circuit-breaking directly controls the representations that are responsible for harmful outputs in the first place. Our technique can be applied to both text-only and multimodal language models to prevent the generation of harmful outputs without sacrificing utility—even in the presence of powerful unseen attacks. Notably, while adversarial robustness in standalone image recognition remains an open challenge, circuit breakers allow the larger multimodal system to reliably withstand image “hijacks” that aim to produce harmful content. Finally, we extend our approach to AI agents, demonstrating considerable reductions in the rate of harmful actions when they are under attack. Our approach represents a significant step forward in the development of reliable safeguards to harmful behavior and adversarial attacks. Code is available at github.com/GraySwanAI/circuit-breakers.
Work done while at Gray Swan AI. Work done in advising capacity to Gray Swan AI.
Correspondence to: andy@grayswan.ai
@@ 3 Circuit Breaking with Representation Engineering
In this section, we introduce a novel approach aimed at mitigating the generation of harmful outputs in neural networks by inducing a new type of phenomenon called “circuit-breaking.” This phenomenon can be elicited using a family of techniques designed to monitor or remap model representations related to harmful processes, redirecting them towards incoherent or refusal representations. This process is reminiscent of “short-circuiting,” where harmful representations are “shorted” and intercepted by circuit breakers. The core objective of this method is to robustly prevent the model from producing harmful or undesirable behaviors by through monitoring or controlling the representations.
@@ 4.1 Large Language Models
    Input Embedding Attack: This white-box attack operates in the embedding space by optimizing a set of input embeddings directly instead of using hard tokens, with the objective of eliciting an affirmative assistant response [60].
@@ 4.2 Multimodal Models
Results.  Figure 3 demonstrates that for multimodal models, our circuit-breaking technique RR is also able to make a model significantly more robust while preserving model capabilities. Especially when subject to white-box PGD Attack, RR achieves reduction of $84\%$ in the compliance rate compared to the original model and $85\%$ compared to the safety prompt. Meanwhile, performance on MMMU and LLaVA-Wild remains within $0.5\%$ of the original, as opposed to the safety prompt which causes a decrease of $3.3\%$ on LLaVA-Wild. This demonstrates that despite the ongoing challenge of achieving adversarial robustness in standalone image recognition, circuit breakers enable the larger multimodal system to reliably counter image “hijacks” [6] intended to elicit harmful outputs.
@@ 4.4 Ablation and Analysis
Table 2: Comparison of Harmfulness Probing (HP) and Representation Rerouting (RR). RR is a representation control method whereas HP is a representation reading method. HP, when applied using a reasonable threshold, significantly lowers the Attack Success Rate compared to a refusal-trained baseline but lags behind the performance achieved by RR.
  Mistral-7B-Instruct-v2 Llama-3-8B-Instruct
Refusal Trained + HP + RR Refusal Trained + HP + RR
  Over-Refusal WildChat 2.0 3.6 3.4 2.2 6.2 6.2
  Robustness No Attack 57.8 17.9 4.9 12.4 7.8 1.2
Manual 77.4 7.4 6.8 8.3 2.1 0.0
TAP-T 85.8 27.5 17.5 17.4 6.6 2.1
GCG 88.7 18.0 11.2 44.5 10.0 2.5
Input Embed 92.1 16.3 15.7 80.4 13.7 9.6
Average 77.4 17.7 10.1 20.7 6.6 1.5
Circuit breaking with Harmfulness Probes (HP).  Our proposed circuit breaking method relies on representation control. This section evaluates the potential efficacy of representation reading as an alternative. Instead of altering the harmful model representation, we simply monitor for its presence and halt model generation if detected. We employ the same training dataset (harmful circuit breaker set and retain set) used in the LLM experiments. A logistic regression classifier is trained to distinguish between model activations from the two datasets. Specifically, activations are collected from the 16th layer of the Mistral model and from the final layer of the Llama-3 model for each token in 2000 responses. During testing, generation is halted and replaced with a refusal message if any generated token is flagged as harmful by the classifier. We find a threshold so that the false positive rates (FPR) on WildChat (Appendix B) is around the same as the model trained with RR [74]. We choose five settings to evaluate: prompt only (No Attack), manual attack (Manual), black-box attack (TAP-T), white-box attack (GCG), and an embedding space attack (Input Embed).
As shown in table Table 2, Harmfulness Probing (HP) significantly reduces the attack success rate compared to the refusal-trained baseline, but is outperformed by the representation control approach (RR). We emphasize that, although generally probing can be easily thwarted by adversarial attacks, much like input and output filters, its robustness in this context can be largely attributed to the continuous monitoring of model representations associated with harmful processes throughout the entire generation, a key idea in circuit breaking. It is important to note, however, that the Harmful Probes are tested under a weaker adversarial setting, where the attacker lacks knowledge of the probe and does not directly optimize against it. Further investigation into Harmfulness Probes is left for future work.
@@ 5 Limitations and Conclusion
We are thankful to Steven Basart, Stephen Casper, David Dalrymple, Xander Davies, and Fabien Roger for providing valuable feedback on the paper.